Charles Miller invested his textbook royalty income in two accounts, one paying 3% annual simple interest and the other paying 2% interest. He earned a total of $11 interest. If he invested three times as much in the 3% account as he did in the 2% account, how much did he invest at each rate?
Answer by ptaylor(2198)

Interest (I)=Principal(P) times Rate(R) times Time(T) or I=PRT
Let P=amount invested at 2%
Then 3P=amount invested at 3%
Interest at 2% for 1 year=P*0.02*1
Interest at 3% for 1 year=3P*0.03*1
And we are told that the sum of the above interests =$11. So our equation to solve is:
0.02P+3P*0.03=$11 simplify
0.02P+0.09P=$11 collect like terms
0.11P=$11 divide both sides by 0.11
P=$100-----------------------amount invested at 2%
3P=3*$100=$300-----------------------------amount invested at 3%
CK
100*0.02+300*0.03=11
2+9=11
11=11

Hope this helps---ptaylor